A bin contains 25 lightbulbs, 5 of which are in good condition and will function for at least 30 days, 10 of which are partially defective and will fail in their second day of use, 10 will not light up. Given that a randomly chosen bulb initially lights, what is the probability it will still be working after one week?
I was thinking that the prob. of picking a lightbulb that initially lights is 15/25. And the prob. that it will work more than a week is 15/25 * 5/25. Am i wrong thinking this way?

Answer by sudhanshu_kmr(1152)


Bulb will still be working after one week if it is one of 5 good bulb.
probability = 5/25 = 1/5
yes, thinking that way is wrong.

you are welcome to get online free help: sudhanshu.cochin@gmail.com